A firefighter stands 50m from a building that caught fire at a particular height. He starts shooting water to remove the blaze. The water moves at a speed of 40m/s and the pipe is inclined or projected at 30 degrees above the horizontal. Calculate the height where the fire caught?


1
Expert's answer
2021-08-03T11:42:45-0400

"v = 40\frac{m}{s}"

"s = 50 m"

"\alpha = 30\degree"

"\vec g = 9.8 \frac{m}{s^2}"

"\text{Consider the movement of a stream of water as}"

"\text{a body thrown at an angle to the horizon:}"

"s(t) = v_xt;"

"v_x= v\cos\alpha" ;

"t = \frac{s(t)}{ v\cos\alpha}= \frac{50}{ 40\cos30\degree}=1.44s"


"h(t) = v_yt-\frac{gt^2}{2}"

"v_y= v\sin\alpha"

"h(t) = v\sin\alpha *t-\frac{gt^2}{2}=40\sin30\degree=9.83m"


"\text{Answer: } h = 9.83m"